# Dynamic Context Adaptation for Consistent Role-Playing Agents with Retrieval-Augmented Generations

### v1

@@ Introduction
While RAG has demonstrated significant potential in improving factual correctness of LLMs, existing RAG methods tend to generate uninformative responses (e.g., I’m sorry, but I don’t have specific information.) for questions outside their knowledge base (Guo et al. 2024; Shukla et al. 2025; Wang, Leung, and Shen 2025), or excessively and repetitively use irrelevant chunks that are not pertinent to the given query.
In this paper, GS focuses on selecting appropriate chunks to generate natural and persona-consistent responses. GS is composed of three stages. First, we iterate over the chunks, which are sorted in descending order of similarity to the user query $u$, and employ an LLM to determine whether it is possible to infer the corresponding character’s attributes from each chunk for the $u$. Second, the chunks selected in the previous step are added to the slot, and the iteration terminates when the slot is full. Finally, if the slot remains empty after the maximum number of search iterations, the $K$ chunks with the highest similarity to the query is returned.
GS is effective in identifying chunks containing information that can be inferred from a character’s actions, such as beliefs or personality traits, which are not explicitly stated in the knowledge base and therefore are difficult to retrieve through direct search. For example, if there is no explicit knowledge corresponding to the query ”My living and work spaces are clean and organized” but the information ”Megumin is diligent.” is available, RPA can use this information to infer the characteristics of Megumin and generate an appropriate response. GS is summarized in Algorithm 1.
Attribute Extractor
Inspired by the observations that incorporating character’s attributes can lead to more realistic responses (Park, Park, and Lim 2025; Chen et al. 2025), AE considers two attributes: Belief and Value, and Psychological Traits. Beliefs and values are fundamental principles and ideological orientations that inform and influence a character’s viewpoints and choices. On the other hand, psychological traits refer to characteristics related to personality, emotional states, personal interests, and cognitive tendencies. AE extracts attributes of a character from the chunks generated as a result of GS, and dynamically exploits them as context.
Experiments
Setup
Dataset. We construct the CharacterRAG dataset, which consists of 15 fictional characters, to leverage a RAG-based role-playing framework. CharacterRAG is a high-quality, role-playing dataset in which all external information about works featuring characters that could affect persona consistency has been manually removed, and each persona document has been directly reconstructed from the perspective of each character by human annotators [^1] CharacterRAG dataset is sourced from Namuwiki and is based on Korean data: https://namu.wiki/. For instance, any information speculated from the perspective of editors rather than the characters themselves, as well as information such as character popularity polls that may disrupt role-playing, is excluded. CharacterRAG consists of 15 distinct fictional characters, 976K written char- acters, and 450 question–answer pairs. Six attributes [^2] Unlike the other attributes, Belief and Value and Psychological Traits directly impact the behavior of a character; therefore, AE utilizes information from these two attributes. define each character’s persona and the corresponding QA pairs (Chen et al. 2025):
  • Activity: A documented history comprising prior activities, behaviors, and interactions, encompassing elements such as backstory and schedules.
  • Belief and Value: The foundational tenets, dispositions, and ideological orientations that inform and guide a character’s viewpoints and decision-making processes (e.g., beliefs and attitudes).
  • Demographic Information: Information that can identify a character, including but not limited to their name, age, educational background, professional history, and geographic location.
  • Psychological Traits: Attributes associated with personality traits, emotional states, preferences, and patterns of cognitive behavior.
  • Skill and Expertise: The extent of understanding, skillfulness, and competence regarding particular fields or technologies.
  • Social Relationships: The characteristics and processes of social interactions, encompassing individuals’ roles, relational ties, and patterns of communication.
[…]
$\sum\mu$ $\sum\sigma^{2}$ $\sum\mu$ $\sum\sigma^{2}$ $\sum\mu$ $\sum\sigma^{2}$
Vanilla 6.4325 0.1026 8.3306 0.1557 12.3136 0.0071
ATS 6.7007 0.0884 8.4718 0.1281 12.2336 0.0070
ACTS (Ours) 6.8575 0.0784 8.6226 0.1179 12.3240 0.0063
Table 1: Distribution of Similarity Scores. Vanilla refers to the method of splitting chunks with a fixed length and overlap, ATS refers to dividing a persona for each character with optimal segment length and overlap, without hierarchical context.
$\mu$ $\sigma$ Mdn Cronbach’s alpha
BFI 3.890 0.965 4.100 0.845
MBTI 3.870 0.911 3.933 0.830
Table 2: Human Evaluation. We conduct human evaluation using a 5-point Likert scale to verify whether inferring character attributes with the Attribute Extractor (AE) from chunks extracted by Guided Selection (GS) is reasonable. Note that $S$ represents the results of all human evaluators, $\mu$ is $\mathbb{E}(\mathbb{E}(S))$, $\sigma$ is $\mathbb{E}(\sigma(S))$, and Mdn is $\mathbb{E}(Mdn(S)).$
Chunk Usage Rate GS Failure
CharacterRAG 67.56% 25.11%
MBTI 43.84% 15.33%
Table 3: Average Chunk Usage Rate and GS Failure. We present 30 CharacterRAG questions and 60 MBTI questions to 15 characters, and measure the average chunk usage rate and GS failure.
Each section of the character’s persona contains subsections, preserving the hierarchical context of each chunk (e.g., ”Frieren’s activities in the story, the journey to find Aureole, the First-Class Mage Exam”). Furthermore, Each QA pair consists of a question and corresponding answer derived from the character’s knowledge, pertaining to one of the six attributes manually constructed for each character. Detailed statistics of CharacterRAG are shown in Figure 3.
[…]
Settings. We implement Guided Selection (GS) and Attribute Extractor (AE) using GPT-4.1 (”gpt-4.1-2025-04-14”). We apply the ACTS to the conventional Naive RAG. The maximum number of search iterations $N$ is 30, and the slot size $K$ is set to 2. Note that the speed at which GS identifies and retrieves relevant chunks for answering is much faster and more cost-efficient than that of previous graph-based or web search-based RAG methods (Shukla et al. 2025; Guo et al. 2024; Yan et al. 2024).
Evaluation Protocols
Tasks. We use 450 QA pairs from the CharacterRAG dataset to verify whether the RPA sufficiently leverages each character’s knowledge. We also use 60 MBTI [^3] https://www.16personalities.com/ questions and 120 BFI (Barrick and Mount 1991) questions to investigate whether each character can appropriately respond to questions for which they do not have explicit prior knowledge. Since it is not possible to construct QA pairs for questions outside the scope of a character’s knowledge, we instead conduct an interview-based assessment (Wang et al. 2024b) for each character and compare the results to psychological test outcomes for the character, as determined by thousands of actual participants’ votes [^4] https://www.personality-database.com/.
Metrics. We design three metrics to extensively evaluate the role-playing capabilities of RAG-based RPAs: (i) ACC measures whether the character’s response contains the correct answer or not. (ii) ACCL is a score assigned by the LLM, ranging from 1 to 10, that evaluates how well the character’s response reflects the correct answer. (iii) Hallucination Score (HS) evaluates the degree of hallucination in the model’s response given a query, the relevant chunks, and the ground-truth answer, on a scale from 1 to 10. Specifically, HS is assigned close to 1 when the response faithfully reflects only the facts contained in the chunks or answer without distortion or addition, indicating minimal hallucination.
CRAG LightRAG AMADEUS (Ours) GT
Anya Forger INTJ (-3) INFJ (-2) ENFP (0) ENFP
Chika Fujiwara ENFP (0) INTP (-2) ENFP (0) ENFP
Edward Elric ISTJ (-3) INTP (-1) INFP (-2) ENTP
Frieren INFP (-1) INTP (0) INTP (0) INTP
Hitori Gotoh ISTJ (-2) ENFP (-1) INFP (0) INFP
Light Yagami INTJ (-1) INTJ (-1) INTJ (-1) ENTJ
Māo Māo INTJ (-1) INTJ (-1) ISTP (-1) INTP
Megumin INFP (0) INFP (0) ISFP (-1) INFP
Mikoto Misaka ISFP (-4) ENTJ (0) INFJ (-2) ENTJ
Nina Iseri ISFP (0) ENFJ (-3) INFP (-1) ISFP
Saitama ISTP (0) INTP (-1) ISTP (0) ISTP
Son Goku ENFJ (-2) INTP (-3) ESFP (0) ESFP
Tanjiro Kamado ENFJ (0) INTP (-3) ENFJ (0) ENFJ
Tobio Kageyama ENTJ(-2) INFJ (-1) ISTJ (0) ISTJ
Yui Hirasawa ENFP (0) INTP (-2) ESFP (-1) ENFP
Accuracy 68.33% 65.00% 85.00% -
Avg F1-Score 0.6448 0.5344 0.8244 -
Table 4: Predicted MBTI Types per Character. The number in parentheses indicates the number of times the ground-truth (GT) type of each character was not correctly identified (using the GPT-4.1 setting).
Experimental Results
[…]
CRAG LightRAG AMADEUS (Ours) GT
Anya Forger SLOAI (-2) RCUEN (-3) SLUEI (-2) SCUAI
Chika Fujiwara SCUAI (0) RLUEN (-4) SCOAI (-1) SCUAI
Edward Elric SLOEI (-1) RCUAN (-4) SLOEI (-1) SLUEI
Frieren RCUAI (-1) SLUEN (-3) RCUAI (-1) RCUEI
Hitori Gotoh RLUAI (0) RCUEN (-3) RLUAI (0) RLUAI
Light Yagami SCOEI (-1) RCUAN (-3) SCOEI (-1) RCOEI
Māo Māo RCOAN (-2) SLUAN (-5) RCOEI (0) RCOEI
Megumin SCUAI (-2) RLUEN (-2) SLOEI (-1) SLUEI
Mikoto Misaka SLOAI (-3) RCUEN (-2) SLOAI (-3) RCOEI
Nina Iseri RLUAI (-1) RCUEN (-2) SLUEI (-1) RLUEI
Saitama RCOAN (-1) SCOAI (-3) RCUAN (0) RCUAN
Son Goku SCUAI (-1) RLUEI (-4) SCOAI (-2) SCUAN
Tanjiro Kamado SCOAI (0) RCUAN (-3) SCOAI (0) SCOAI
Tobio Kageyama SLOAN (-2) RCUAI (-4) RCOEN (-1) RLOEN
Yui Hirasawa SLUAI (-1) RCOEN (-4) SCUAI (0) SCUAI
Accuracy 76.00% 34.67% 81.33% -
Avg F1-Score 0.7313 0.2774 0.7986 -
Table 5: Predicted Big 5 SLOAN Types per Character. The number in parentheses indicates the number of times the ground-truth (GT) type of each character was not correctly identified (using the GPT-4.1 setting).
Extracting a character’s attributes from selected text chunks is reliable. We investigate the reasonableness of inferring character attributes with the Attribute Extractor (AE) from chunks extracted via Guided Selection (GS) by conducting human evaluation using a 5-point Likert scale. To this end, we invite 11 human evaluators and each evaluator is asked to score 60 randomly selected samples. Each sample consists of pairs of chunks selected from GS and attributes extracted through AE, for 30 BFI questions and 30 MBTI questions that are not included in the character’s knowledge.
In Table 2, we find that the means $\mu$ is close to 4, with small standard deviations $\sigma$. It demonstrates that the outputs of GS and AE are reliable and trustworthy, even from a human evaluative perspective. We also measure Cronbach’s alpha (Cronbach 1951) to evaluate internal consistency among the human evaluators. We find that the Cronbach’s alpha values are 0.845 and 0.830, both exceeding the commonly accepted threshold of 0.7 for acceptable reliability. Since values above 0.8 are generally interpreted as indicating a high level of internal consistency, experimental results in Table 2 can be considered highly trustworthy.
Selecting appropriate chunks increases the utilization of a character’s persona. Table 3 shows the usage rate of persona chunks and the failure rate of GS when each each of the 15 characters is presented with questions that are both within and outside their knowledge. Note that, despite the limited number of queries in CharacterRAG, the Chunk Usage Rate still reaches 67.56%. The findings demonstrate that GS serves as an effective tool for enabling the use of character personas. At the same time, it can be observed that the 450 questions in CharacterRAG are designed to capture the core information of each character.
Meanwhile, as shown in Table 3 and Figure 1, GS substantially increases the chunk utilization rate for the 60 MBTI questions, from 34.93% to 43.84%. The proportion of cases where the GS fails to find an appropriate chunk, i.e., GS Failure, is 25.11% and 15.33%, respectively, which demonstrates that although GS effectively utilizes the character’s persona, there is still room for further improvement. Nevertheless, despite this shortcoming, our framework offers markedly enhanced speed, cost-effectiveness, and stable performance compared to earlier RAG approaches utilizing graph structures or web search mechanisms. (Shukla et al. 2025; Guo et al. 2024; Yan et al. 2024).
Method Task GPT-4.1 Gemma3-27B Qwen3-32B
[…]
w/o RAG CharacterRAG 49.56% 6.79 - 27.56% 5.33 - 18.89% 4.35 -
Naive RAG CharacterRAG 91.33% 9.23 3.13 86.44% 8.85 3.27 78.44% 8.49 5.05
LightRAG CharacterRAG 48.00% 6.06 - 69.56% 8.17 - 68.67% 8.20 -
CRAG CharacterRAG 70.00% 8.26 3.21 57.78% 7.57 4.09 28.67% 5.24 8.68
AMADEUS (Ours) CharacterRAG 92.67% 9.26 2.89 88.00% 8.92 3.26 78.89% 8.63 4.66
Naive RAG MBTI - - 2.69 - - 2.53 - - 2.33
CRAG MBTI - - 2.38 - - 2.91 - - 1.80
AMADEUS (Ours) MBTI - - 2.05 - - 2.02 - - 2.04
Naive RAG BFI - - 2.74 - - 2.52 - - 2.42
CRAG BFI - - 2.26 - - 2.75 - - 1.96
AMADEUS (Ours) BFI - - 1.94 - - 1.99 - - 2.03
Table 6: Evaluation of role-playing capabilities on CharacterRAG, MBTI, and BFI tasks. Higher values of ACC (%) and ACCL (1-10) correspond to better performance, whereas lower values of HS (1-10) are preferable.
Graph-based RAG and web search-based RAG are unsuitable for role-playing. One of the major challenges in retrieval-based role-playing is that, when a RPA receives questions involving knowledge outside a character’s persona, it tends to either overuse irrelevant chunks (Figure 1) or generate uninformative responses (Guo et al. 2024; Shukla et al. 2025; Wang, Leung, and Shen 2025).
To investigate whether RAG-based RPAs can handle this problem, we conduct extensive experiments in which we ask 15 characters 60 MBTI questions and 120 BFI questions each, and and evaluate their ability to accurately infer the characters’ personality types. Table 4 and Table 5 show predicted MBTI types and Big 5 SLOAN types per character. Our framework maintains persona consistency even when answering questions that are not explicitly specified in each character’s persona in both MBTI and BFI settings. Note that the performance gap of CRAG is significant between the two settings. We assume that questions requiring analogical reasoning are difficult to solve even with web search and that the search results may contain non-negligible noise. On the other hand, LightRAG exhibits the lowest performance, which shows that graph-based RAG methods are not well suited for RPA applications due to the high cost of graph construction, the difficulty in adding or removing new knowledge, and challenges in maintaining persona consistency. While we did not perform a direct comparative experiment, we observed that GraphRAG (Shukla et al. 2025) suffers from similar problems. Specifically, applying graph-based RAG to RPA is challenging in cases where the persona frequently changes, such as with daily streamers or VTubers.
CharacterRAG dataset serves as a valuable resource for the construction and evaluation of RAG-based RPAs. To investigate the factors influencing the performance of role-playing, we conduct a comprehensive interview-based assessments on the generalization capabilities of models with various LLMs and RAG techniques. Table 6 presents how the ability to accurately answer questions related to the character’s knowledge, which is a core aspect of role-playing, varies across the applied methodologies.
We first examine to what extent each LLM possesses background knowledge about the 15 characters in a setting without RAG, and results show that none of the three LLMs are capable of effective role-playing without access to external knowledge. Moreover, we observe that LightRAG, a graph-based RAG, is ill-suited for the storage and retrieval of character knowledge, as it often suffers from issues such as entity ambiguity and uninformative responses. In a similar vein, CRAG exhibits challenges in maintaining role-playing fidelity, which can be attributed to the tendency of web search-based RAG methods to utilize retrieved content that may undermine the consistency of a character’s persona. Indeed, despite leveraging web information, CRAG is able to correctly answer only 6 out of the 30 CharacterRAG questions pertaining to Nina Iseri. In addition, to analyze how a thinking mode of LLMs influences their role-playing capabilities, we employ Qwen 3-32B. Results demonstrate that the thinking mode fails to yield any substantial positive effect on enhancing role-playing performance.
Note that our framework achieves the best performance across all three LLMs. We also find that the Hallucination Score (HS) is the lowest in CharacterRAG setting. These results highlight the importance of preserving the context of split-character personas and effectively leveraging appropriate character attributes in RAG-based RPAs. Furthermore, such elements are especially pronounced in dialogue situations that transcend the scope of the character’s knowledge (Table 4 and Table 5). We believe that our findings demonstrate new possibilities for RAG-based RPAs.
Conclusion

### v3

@@ Appendix A Chunk Duplication Frequencies
In Figure 6, previous RAG method retrieves (k = 3) relevant chunks for each of the 60 given MBTI questions, and subsequently measures both the chunk utilization rate and the duplication frequency. Our method measures the chunk utilization rate and duplication frequency for the chunks selected by ACTS and GS.
Appendix B Genre Diversity in CharacterRAG
Table 7 shows the genres of works in which each character appears. Note that while there are some similar categories, CharacterRAG encompasses a broader range of genres. The detailed personality distribution for each character (e.g., Extroverted vs. Introverted) is shown in Figure 2.
Appendix C Related Work
Role‑Playing Agents. With the advent of LLMs, researchers have pursued finer‑grained persona consistency (Zhang et al., 2018; Ji et al., 2025; Park et al., 2025; Lu et al., 2024; Zhou et al., 2024). Complementary benchmarks soon followed, along with various evaluation methods (Boudouri et al., 2025; Ahn et al., 2024; Wang et al., 2024b, 2025a). However, there has been little research on RAG-based role-playing agents (RPAs). In this paper, we propose AMADEUS, a RAG-based RPA framework that not only elicits information related to a character, but also maintains persona consistency even when responding to queries beyond its explicit knowledge.
Retrieval‑Augmented Generation (RAG). RAG couples non-parametric memory with an LLM to mitigate hallucination and stale knowledge (Lewis et al., 2020; Bhat et al., 2025; Zhong et al., 2025b; Zhu et al., 2024). Nevertheless, no existing benchmark explicitly targets role‑playing under RAG, and prior work still assumes personas are short, knowledge‑dense snippets. We mitigate this discrepancy with a novel text splitter that tailors chunk lengths and hierarchical context to each character. We also introduce CharacterRAG, the first dataset designed for building and evaluating RAG‑based role‑playing agents across 15 fictional character’s personas.
Table 8: Role-playing capabilities on MBTI and BFI. Lower values of HS (1-10) are preferable.
RAG Method GPT-4.1 Gemma3-27B Qwen3-32B
HS ($\downarrow$) HS ($\downarrow)$ HS ($\downarrow$)
MBTI
Naive RAG 2.69 2.53 2.33
CRAG 2.38 2.91 1.80
AMADEUS (Ours) 2.05 2.02 2.04
BFI
Naive RAG 2.74 2.52 2.42
CRAG 2.26 2.75 1.96
AMADEUS (Ours) 1.94 1.99 2.03
Prompt used in Guided Selection In the given [persona_name]’s Context, output True if you can clearly infer the Belief and Value information or Psychological Traits of [persona_name] to answer the given User Query; otherwise, output False. Output only True or False.
# Explanation of Belief and Value:
- The principles, attitudes, and ideological stances that shape an individual’s perspectives and decisions
- e.g.,
[persona_name]’s "beliefs", "attitudes", "values", "religion"
# Explanation of Psychological Traits:
- Characteristics related to personality, emotions, interests, and cognitive tendencies.
- e.g.,
[persona_name]’s "Personality", "hobby and interest", "emotional"
Figure 3: Prompt used in Guided Selection (GS). GS returns whether Belief and Value information or Psychological Traits can be inferred from given chunks.
Prompt ued in the interview-based psychological assessment (MBTI) Based on the given [persona_name]’s answer, please rate the question, "Am I ([persona_name]) [question]?" on a scale from 1 to 7.
A score closer to 1 means "very much so," and a score closer to 7 means "not at all."
Output only the number.
Figure 4: Prompt ued in the interview-based psychological assessment (MBTI). In the BFI test, responses are rated on a scale from 1 to 5.
Figure 5: AMADEUS framework. AMADEUS consists of three substage: (i) ACTS splits a persona document to make it suitable for RAG-based role-playing. (ii) To fully leverage the character’s knowledge, GS retrieves chunks from which it can infer the answer to the user query. (iii) AE uses the information derived from the GS results to extract character attributes.
Figure 6: Chunk Duplication Frequencies. We compare the distribution of chunk duplication frequencies and chunk usage rates between Naive RAG and our method when questions involving knowledge not present in the persona document were given. We observe that when each of the 60 MBTI questions is asked to 15 characters, the average chunk usage rate increases from $34.93\%$ to $43.84\%$, and the distribution becomes more uniform.
Figure 7: An example of the output produced at each substage of Amadeous. To improve interpretability, we provide a detailed breakdown of the results for a character from the CharacterRAG dataset (Yui Hirasawa).
Appendix D Settings
We implement GS and AE using GPT-4.1 ("gpt-4.1-2025-04-14"). We develop the Amadeus framework based on naïve RAG. The maximum number of search iterations $N$ is 30, and the slot size $M$ is set to 2. We performed benchmarking on an NVIDIA H100 NVL GPU.
Appendix E Log-Density Ridgeline Analysis
To empirically validate the suitability of the overlap coefficient, Figure 8 presents the log-density ridgelines of five distributions estimated under the normality assumption: $\log f\!\left(x\mid\sum\mu,\,\sum\sigma^{2}\right)=-0.5\left(\frac{x-\sum\mu}{\sqrt{\sum\sigma^{2}}}\right)^{2}-\log\!\left(\sqrt{\sum\sigma^{2}}\right)-0.5\log(2\pi)$. We observe that, when $\alpha=2$, the sum of the similarity scores is maximized while their variance is minimized.
Figure 8: Ridgeline of Log-Density Comparison. Empirical verification of optimal overlap coefficient $\alpha$: $l_{\text{o}}=l_{\text{max}}/\alpha$ (Normal assumption).